a retailer an account for 20,000 naira from a manufacturer who give two months credit ,but offers a cash discount of 5 percent for prompt payment. the retailer borrow 20000 naira from his bank 9 percent per annum for two months and pays promptly to secure the cash discount. how much does he gain by this transaction?


1
Expert's answer
2020-11-26T19:55:29-0500

How much is the a cash discount for prompt payment?


"20000\ naira\times \dfrac{5\%}{100\%}=1000\ naira"

Calculate Simple Interest


"S.I.=20000\ naira\times \dfrac{9\%}{100\%}\times\dfrac{2}{12}=300\ naira"

How much does he gain by this transaction?


"1000\ naira-3000\ naira=700\ naira"

He gains 700 naira by this transaction.